OpenStudy (anonymous):

If x= y-3 and y=z^2, what is x in terms of z

OpenStudy (anonymous):

x in terms of z means you want some equation with z equal to x this can be by just substituting for y into the equation x=y-3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so do you know what to do?

OpenStudy (anonymous):

what does ^ that mean

OpenStudy (anonymous):

\[x^2\]=x^2 it means to the power of

OpenStudy (anonymous):

k so how do i get my answer

OpenStudy (anonymous):

you know that y=z^2 y is equal to z^2 you want to solve x in terms of z you substitute for y x=y-3 x=(z^2)-3 does this make sense

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

and that's just the answer

OpenStudy (anonymous):

oka

OpenStudy (anonymous):

because its an equation with z that equals x

OpenStudy (anonymous):

but the answer choices are different

OpenStudy (anonymous):

answer choices?

OpenStudy (anonymous):

z^-6 b. z^6 c. z^-2/3 d. z^5 e. z^3/2

OpenStudy (anonymous):

-6?

OpenStudy (anonymous):

no idea, @satellite73 @panlac01 did i do something wrong?

OpenStudy (anonymous):

@julialeigh are you sure your question wasnt \[x=y^{-3}\] \[y=z^2\] instead?

OpenStudy (anonymous):

x=y^(-3) y=z^2

OpenStudy (anonymous):

@julialeigh please rewrite the original equation the way you have it written, the answer isnt one of the choices the way you have it written

OpenStudy (anonymous):

lol it must be \(y^{-3}\) right?

OpenStudy (anonymous):

if it is x=y^-3 then substitute y \[x=(z^2)^{-3}\] when you find the power of the an exponent, you multiply the 2 numbers \[x^{a^b}=x^{a*b}\] so \[z^{2^{-3}}=z^{2*-3}\]